OpenStudy (anonymous):

Medalpolooza A right triangle has a hypotenuse of length 10 yd and one leg of length 2 yd. How long is the other leg?

OpenStudy (butterfield1215):

ok for this you will be using the pythagorean thereom which is \[a^2+b^2=c^2\]

OpenStudy (anonymous):

ok :)

OpenStudy (butterfield1215):

so 2*2 is 4 and 10*10 is 100 so 4+x=100

OpenStudy (butterfield1215):

subtract 4 from both sides and find the sqare root of it so 96 or 9.7

OpenStudy (butterfield1215):

9.7 squared sorry

OpenStudy (anonymous):

oh ok lol, so after i find out what 9.7 squared is, that will be the answer?

OpenStudy (butterfield1215):

9.7 is your answer

OpenStudy (anonymous):

hmm ok, thats not an answer choice so i guess i have to convert it into something

OpenStudy (anonymous):

but thanks :)

OpenStudy (butterfield1215):

what are the answers then

OpenStudy (butterfield1215):

if 9.8 is one then that will be it

OpenStudy (anonymous):

i dont know where the "radical" button on the keyboard is but ill type them out using the"/" sign as a radical sign

OpenStudy (butterfield1215):

yes its 9.8 i just checked with a pythagorean thereom calculator and it came up 9.8

OpenStudy (butterfield1215):

wait is it a right triangle

OpenStudy (anonymous):

the choices say 4/6 yd 16/6 yd 2/2 yd 8 yd

OpenStudy (anonymous):

yes a right triangle

OpenStudy (anonymous):

yd meaning yards

OpenStudy (butterfield1215):

i would say 8 yds cause that the closest answer to what im getting

OpenStudy (anonymous):

lol ok i was thinking that too for a bit. thanks :)

OpenStudy (butterfield1215):

No problem